OpenStudy (anonymous):

Put in slope-intercept form: 6x+4y=8

hartnn (hartnn):

Hi @Dgirl828 :) \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) To get the slope intercept form, can you try to isolate y from 6x +4y = 8 ?? start by subtracting 6x from both sides.

OpenStudy (anonymous):

okay so that would be 4y=-6x+8?

hartnn (hartnn):

yes! now divideboth sides by 4 :)

OpenStudy (anonymous):

y=-6/4x+2

hartnn (hartnn):

yesss -6/4 can be reduced too.

OpenStudy (anonymous):

-3/2 so it would be y=-3/2x+2!

hartnn (hartnn):

-6/4 = -3/2 so its y = (-3/2) x +2 with m = -3/2 and y-intercept = 2 yes! good ! :)

OpenStudy (anonymous):

Thankyou!

hartnn (hartnn):

welcome ^_^